OpenStudy (anonymous):

divide (32x^3-48x^2-64x-16)/(8x+4)

OpenStudy (anonymous):

8(4x^3-6x^2-64x-16)/4(x+1)

OpenStudy (nikvist):

\[\frac{32x^3-48x^2-64x-16}{8x+4}=\frac{16(2x^3-3x^2-4x-1)}{4(2x+1)}=\]\[=\frac{4(2x^3+x^2-4x^2-2x-2x-1)}{2x+1}=\]\[=\frac{4(x^2(2x+1)-2x(2x+1)-(2x+1))}{2x+1}=4(x^2-2x-1)\]